What is the value of x in the equation 6(4x+5)=3(x+8)+3
Answer by stanbon(75887)

What is the value of x in the equation:
6(4x+5)=3(x+8)+3
-------------
Divide thru by 3 to get:
2(4x+5) = (x+8) + 1
-------------
8x+10 = x+9
7x = -1
x = -1/7
=================
Cheers,
Stan H.
==================